Help pls!!!!!!! A man spends one-third of his money and $ 50 more. After that, one-fifth of his money and $ 350 more was left to him. How much did he have at first?
Answer by rothauserc(4718)

let x be how much money he had at first, then
x/3 + 50 = x/5 + 350
subtract 50 from both sides of =
x/3 = x/5 + 300
subtract x/5 from both sides of =
5x/15 - 3x/15 = 300
2x/15 = 300
2x = 4500
x = 2250
at first he had $2250